OpenStudy (anonymous):

Revenue, Cost, And Profit Functions? Fees $4 Per Novel $900 Per Month Operating Cost $5.50 per Novel

OpenStudy (anonymous):

R=5.50n C=900+4n P=1.50n-900 n=numeber of novels sold/produced